One number is 4 more than another. If 6 times the smaller minus 4 times the larger is 4, what are the two numbers?
Second question the length of a rectangle is 5in more than twice the width. If the perimeter of the rectangle is 76 in what are the two dimensions of the rectangle?
I understand the examples of similar problems and why the answers are correct but when I do them I'm placing the equations the wrong way and not understanding why. Would greatly appreciate any help in this matter.

Answer by stanbon(48546)

One number is 4 more than another. If 6 times the smaller minus 4 times the larger is 4, what are the two numbers?
Let the smaller number be "x"; The larger is "x+4"
EQUATION:
6x-4(x+4)=4
2x-16=4
2x = 20
x = 10 (smaller number)
x+4=14 (larger number)
-----------------------------
Second question the length of a rectangle is 5" more than twice the width. If the perimeter of the rectangle is 76 in what are the two dimensions of the rectangle?
---------------
Let the width be "x" inches; Length = "2x+5" inches.
EQUATION:
Perimeter = 2(width + length)
76 = 2(x + 2x+5)
38 = 3x+5
3x = 33
x = 11 inches (width)
2x+5 = 27 inches (length)
=============
Cheers,
Stan H.